How do you do the following?
x2-7x+10=0
x2+5x+6=0
x2-3x-4+0
y2+2y-8=0

Answer by checkley75(3666)

X^2-7X+10=0
(X-5)(X-2)=0
X-5=0
X=5 ANSWER.
X-2=0
X=2 ANSWER.
-----------------------------------------------
X^2+5+6=0
(X+3)(X+2)=0
X+3=0
X=-3 ANSWER
X+2=0
X=-2 ANSWER
----------------------------------------------
I DID 2 NOW YOU TRY THE NEXT 2!!!!!